CH4 + 2O2 ----> CO2 + 2H2O  Heat released is 890.4kJ


How much heat is released by burning 15g of CH4?


1
Expert's answer
2021-04-16T05:10:23-0400

Solution:

The balanced thermochemical equation:

CH4 + 2O2 → CO2 + 2H2O, △H = -890.4 kJ

 

The equivalences for this thermochemical equation are:

1 mol CH4 ⇔ 2 mol O2 ⇔ 1 mol CO2 ⇔ 2 mol H2O ⇔ -890.4 kJ

 

Convert the amount of methane (CH4) to moles:

15 g CH4 × (1 mol CH4 / 16 g CH4) = 0.9375 mol CH4

 

Use the thermochemical equation to determine the heat:

0.9375 mol CH4 × (890.4 kJ / 1 mol CH4) = 834.75 kJ

 

Answer: 834.75 kJ of heat is released by burning 15 g of CH4.